the maximum current that can be measured by a galvanometer of resistance 40 ohm is 10mA. It is converted into a voltmeter that can read upto 50V. the resistance to be connected in series with the galvanometer is

Grade:12th pass

3 Answers

Arun
5 years ago
To convert a galvanometer into voltmeter, the necessary value of resistance to be connected in series with the galvanometer is
R = V/ Ig  -  G
R = 5000 - 40
R = 4960 ohm
Gitanjali Rout
5 years ago
To convert a galvanometer into voltmeter, the necessary value of resistance to be connected in series with the galvanometer isR = V/ Ig  -  GR = 5000 - 40R = 4960 ohm
thanks
 
Gitanjali Rout
5 years ago
To convert a galvanometer into voltmeter, the necessary value of resistance to be connected in series with the galvanometer isR = V/ Ig  -  GR = 5000 - 40R = 4960 ohm.